OpenStudy (anonymous):

Sonia wishes to make a deposit.She has these bills:10 ones,4 fives,5 tens,and 2 twenties.She has 7 nickels,14 dimes,and 19 quarters.Sonia has checks for $42 and $38.50.Sonia wants to receive $20 in cash.What should she enter on the NET Deposit line?

OpenStudy (anonymous):

total value of bill=10*1+4*5+5*10+2*20=?+20=? Total cash=42+38.50=? find the difference=?